Compute gross profit earned by the company for each of the four costing methods. For specific identification, the March 9 sale consisted of 100 units from beginning inventory and 230 units from the March 5 purchase; the March 29 sale consisted of 80 units from the March 18 purchase and 120 units from the March 25 purchase.hapter 05 Pre But Problems X C Get Homework Help With Chego X G Hemming uses a perpetual inver X OX Sign out @ sources - Googhapter 05 Pre But Problems X C Get Homework Help With Chego X G Hemming uses a perpetual inver X OX Sign out @ sources - Goog
Answer #1

Warnerwoods Company

FIFO LIFO Average cost Specific identification
Sales (330 units × $ 87.40 per unit)+(200 units × $ 97.40) $ 48,322 $ 48,322 $ 48,322 $ 48,322
Less: Cost of goods sold

{(170 units × $ 52.4 per unit)+(160 units × $ 57.4 per unit)} + {(100 units × $ 62.4 per unit)+(100 units × $ 57.4 per unit)}

= (18,093+11,980)

= ($ 30,073)

{(260 units × $ 57.4 per unit)+(70 units × $ 52.4 per unit)}+{(100 units × $ 52.4 per unit)+(100 units × $ 62.4 per unit)}

= (18,529+11,480)

= ($ 30,009)

(530 units × $ 59.08 per unit)

=($ 31,312)

{(100 units × $ 52.4 per unit)+(230 units × $ 57.40 per unit)} + {(80 units × $ 62.4 per unit)+(120 units × $ 64.4 per unit)}

= (18,442+12,720)

= ($31,162)

Gross Profit

= 48,322-30,073

= $ 18,249

= 48,322-30,009

= $ 18,313

= 48,322-31,312

$ 17,010

= 48,322-31,162

= $ 17,160

Average cost = total cost of available units / available units

= {(170 units × $ 52.40 per unit)+(260 units × $ 57.40 per unit)+(120 units × $ 62.40 per unit)+(220 units × $ 64.4 per unit)} / 770 Units

= (8,908+14,924+7,488+14,168)/770 units

= $ 45,488/770 units = $ 59.08 per unit